# NeRC: Neural Ranging Correction through Differentiable Moving Horizon Location Estimation

### v1

@@ Abstract.
GNSS localization using everyday mobile devices is challenging in urban environments, as ranging errors caused by the complex propagation of satellite signals and low-quality onboard GNSS hardware are blamed for undermining positioning accuracy. Researchers have pinned their hopes on data-driven methods to regress such ranging errors from raw measurements. However, the grueling annotation of ranging errors impedes their pace. This paper presents a robust end-to-end Neural Ranging Correction (NeRC) framework, where localization-related metrics serve as the task objective for training the neural modules. Instead of seeking impractical ranging error labels, we train the neural network using ground-truth locations that are relatively easy to obtain. This functionality is supported by differentiable moving horizon location estimation (MHE) that handles a horizon of measurements for positioning and backpropagates the gradients for training. Even better, as a blessing of end-to-end learning, we propose a new training paradigm using Euclidean Distance Field (EDF) cost maps, which alleviates the demands on labeled locations. We evaluate the proposed NeRC on public benchmarks and our collected datasets, demonstrating its distinguished improvement in positioning accuracy. We also deploy NeRC on the edge to verify its real-time performance for mobile devices.
@@ 1. Introduction
This picture shows that the multipath/NLOS effect is related to the relationship among users, satellites, and the environment.
@@
This picture indicates three types of learning paradigms for the regression of pseudorange errors
The ranging errors caused by multipath and NLOS effects are strongly dependent on the complex relationship among satellites (Tx), receivers (Rx), and the surroundings (Sun et al., 2023; Weng et al., 2024b; Sun et al., 2025). Although modeling errors in ranging measurements is challenging, they often exhibit a periodical pattern at a given site because GNSS satellites complete an orbital cycle approximately every 12 hours (Axelrad et al., 2005; Phan et al., 2013). 1(b) shows the periodic ranging errors observed in GNSS data collected with an Android phone at the same site and time across multiple days. The color of each satellite’s trajectory indicates the value of its ranging error, estimated from ground-truth user locations (Weng et al., 2024b) after removing all known error sources, including satellite clock offsets, atmospheric delays, and relativistic effects (Kaplan and Hegarty, 2017).
The predictability of ranging errors has propelled the community away from traditional model-based approaches (Axelrad et al., 2005; Liu et al., 2018b; Bai et al., 2020) toward data-driven methods (Sun et al., 2020, 2021; Zhang et al., 2021; Sun et al., 2023; Weng et al., 2024b; Sun et al., 2025). However, obtaining accurate labels for ranging errors—critical for training neural models—is both difficult and impractical. As illustrated by Figure 2a, previous studies have used various sophisticated algorithms to estimate these errors based on ground truth locations provided by high-performance specialized positioning systems (Zhang et al., 2021; Sun et al., 2023; Weng et al., 2024b; Sun et al., 2025). With the derived ranging errors, an intermediate loss can be computed to train a neural network for ranging correction. Yet, manually calculating ranging errors is inherently imprecise due to the unobservable user clock offsets and stochastic noise (Weng et al., 2024b). This raises the question: can we instead utilize these location labels directly, given that our ultimate goal is localization? If so, we could bypass the need to compute intermediate losses using ranging error labels and instead train neural modules end-to-end with the final positioning task loss.
@@
Can we integrate the two learning paradigms (shown in Figure 2a and 2b) to harness the powerful regression capabilities of neural networks while preserving our interpretable knowledge of GNSS localization? The key ingredient to enable such an end-to-end learning framework should be a differentiable positioning engine, as illustrated in Figure 2c. This engine leverages the outputs from upstream neural models to compute locations while backpropagating gradients from the final positioning task to refine the neural models. This structured learning paradigm, widely adopted in fields such as robotics (Wang et al., 2024; Yang et al., 2023a; Wang et al., 2025), autonomous driving (Huang et al., 2024), and 3D vision (Mildenhall et al., 2020; Chen et al., 2022), has recently been extended to GNSS positioning (Xu et al., 2023a; Weng et al., 2024c; Xu and Hsu, 2024; Hu et al., 2025).
Among the most relevant studies, E2E-PrNet and pyrtklib train upstream neural network to predict pseudorange errors through a differentiable Weighted Least Squares (WLS) localization engine (Weng et al., 2024c; Hu et al., 2025). While these methods improve positioning in rural (Weng et al., 2024c) or lightly urban areas (Hu et al., 2025), they struggle in deep urban environments, as shown in Figure 3. The main reason is that the WLS engine lacks robustness to the severe noise present in such settings, where multipath and NLOS propagation typically cause low carrier-to-noise density ratios () (Weng and Ling, 2023). Furthermore, these methods rely solely on fully labeled 3D location data, limiting their practical deployment since obtaining accurate ground-truth locations is both challenging and costly (Fu et al., 2020).
The evolution of end-to-end approaches for GNSS positioning
@@ 2. NeRC Design
NeRC is a data-driven framework with well-established physical and geometric priors thrown in for ballast. It handles a horizon of raw measurements for neural ranging error regression and robust location estimation. Within it, the neural module is trained with positioning-related metrics. This end-to-end learning pipeline – from raw measurements to locations – comprises two core modules:
@@
NeRC is a sequence-to-sequence framework with raw GNSS measurements as inputs and localization results as outputs. A location-related loss can be computed as a result.
@@ 2.1. Upstream Network for Ranging Correction
Figure 4 illustrates that one input tensor sample has two dimensions corresponding to satellites and features. The MLP operates in a satellite-wise manner—analogous to the position-wise MLPs in the transformer architecture—processing information from all satellites in parallel. A visibility mask handles the time-varying set of visible satellites by zeroing the outputs of invisible ones, analogous to the transformer’s masked softmax. For clarity, Figure 4 presents an unfolded view of the input tensors from time to time . In reality, the NeRC pipeline uses a single MLP that processes the entire time window in parallel. Including the time dimension and the batch dimension for batched training, the resulting four-dimensional input tensor has the shape , where is the batch size, is the total number of satellites, and is the number of the input features. Thus, the MLP can be represented by
where denotes the input horizon of size at time , is the corresponding pseudorange error predictions, and represents all learnable parameters in the neural network.
@@ 2.2.2. Moving Horizon Estimation
where denotes an estimate of an unknown at time given measurements up to time . and represent the covariance matrices of process and measurement noise, respectively. The inverse of them is used to weight the state transition and measurement costs. is the covariance matrix of the state estimation , representing prior knowledge of our confidence in the initial user state for the current horizon . It also represents the “inertia” of the system—a larger increases the contribution of the prior state cost in the overall objective, thereby drawing the state estimates in the current horizon more consistent with historical trajectories. Assuming the system statistics in (7) are Gaussian, MHE formulated by (7) is equivalent to the Maximum A Posteriori (MAP) estimation (Jazwinski, 1970; Weng et al., 2025):
This figure describes the principle of moving horizon estimation.
@@ 2.2.3. Differentiable Formulation of MHE
where is considered as prior knowledge and is initiated using a diagonal matrix with all elements on its diagonal equal to 0.05. is the Jacobian matrix of the nonlinear function :
@@
Figure 5 illustrates the principle of the differentiable MHE. By substituting (9), (10), (11), (12), and (13) into (7), we obtain the state transition cost, measurement cost, and prior state cost. The objective layer aggregates the three types of costs weighted by their corresponding covariance matrices, which is then solved by a general Differentiable Nonlinear Least Squares (DNLS) solver. In our implementation, we employ the Theseus library built for Pytorch (Pineda et al., 2022). The optimizer is initialized with an approximate trajectory guess, denoted by , for each horizon. We leverage the historical state estimates as the approximate trajectory (Weng et al., 2025). The lower half of this diagram depicts the updating process of the prior state covariance matrix using (14), where only the latest covariance matrix gets updated at time . The updated horizon of covariance matrices serves the next horizons.
@@
These figures profile the forward inference and backward training performance of MHE.
@@ 3.2. Forward profiling
6(a) indicates that MHE-w/-AC maintains a lead over other algorithms as the horizon size is small. Unlike the filtering-based MHE-F, which executes a single-pass update at each step, MHE-w/-AC adopts an optimization-based approach that iteratively processes the measurements at every time step. Although MHE-w/o-AC also follows an optimization-based strategy, it only leverages measurements within the current horizon, disregarding prior information. This omission significantly degrades its performance in short horizons. Increasing the horizon size can narrow down the performance gap between MHE-w/-AC and MHE-w/o-AC. Once the horizon extends beyond a certain threshold—approximately 65 in this example—MHE-w/o-AC begins to outperform MHE-w/-AC. At that point, the window is sufficiently long to retain informative measurements for current-state estimation, while older data outside the window degrades performance (Ling and Lim, 1996, 1999; Weng et al., 2025). Nevertheless, continuously increasing the horizon size does not always lead to improved performance. A typical performance curve of MHE-w/o-AC initially decreases to a minimum, then gradually converges back toward that of MHE-w/-AC (Rao and Rawlings, 2002; Weng et al., 2025). In this dataset, we observe that the turning point occurs at a horizon length of approximately 450. The filtering-based MHE-F exhibits the same performance regardless of the chosen horizon size (Muske et al., 1993; Weng et al., 2025).
Regarding the computation time of each method, 6(b) shows that the forward time increases as the horizon size enlarges. Among the methods, MHE-F demonstrates the fastest performance, benefiting from its recursive filtering structure. In contrast, the optimization-based approaches are naturally slower, as they solve a nonlinear least squares problem at each time step. Notably, MHE-w/-AC incurs slightly more computation than MHE-w/o-AC for the same horizon size, due to the additional optimization involved in incorporating the arrival cost. However, MHE-w/-AC can achieve superior positioning accuracy with a smaller horizon, thereby taking less computational time to reach similar performance levels.
Summary: MHE offers advanced positioning capabilities at a competitive cost, making it an attractive choice for downstream positioning engines. Although increasing the horizon size can further enhance the accuracy of MHE-w/o-AC, the marginal gains over MHE-w/-AC with a smaller horizon come with a substantial rise in computational time—a 2.35% improvement in accuracy comes at the expense of approximately 2.4 times the computation time, as shown in 6(a) and 6(b). This trade-off is critical to consider when deploying MHE on resource-constrained platforms such as mobile devices. Additional improvement in MHE is outside the scope of this paper, but can be expected by approximating the arrival cost more accurately (Rao and Rawlings, 2000; Tenny and Rawlings, 2002; Rao et al., 2003; Fiedler et al., 2020) and incorporating constraints on the system (Rao et al., 2001; Haseltine and Rawlings, 2005; Wen et al., 2021; Suzuki, 2023; Xia et al., 2024).
@@ 3.3. Backward profiling
This section investigates the training convergence of the upstream MLP equipped with MHE. 6(c) gives the training loss (3D location loss) after one training epoch for various horizon sizes.
Increasing the horizon size improves training convergence across all estimators. While MHE-w/o-AC exhibits poor convergence at small horizons, it surpasses the other two methods in convergence speed once the horizon exceeds 20, as illustrated in 6(c). It is the arrival cost item in (7)—the system’s ‘inertia’—that decelerates the backward convergence of MHE-F and MHE-w/-AC. Interestingly, it is the same item that enables MHE to achieve satisfactory forward positioning accuracy even with small horizon sizes. Among the two estimators that incorporate the arrival cost, MHE-w/-AC demonstrates better convergence performance than MHE-F, benefiting from its improved positioning accuracy.
6(d) shows the backward training time per batch of the three methods. Just aligned with the forward time, MHE-F exhibits the lowest backward computation time due to its filtering structure. In contrast, both MHE-w/-AC and MHE-w/o-AC exhibit increasingly higher computational costs as the horizon size grows.
Summary: The backward training convergence can be improved by expanding the receding horizon, which is equivalent to reducing the contribution of the arrival cost in the overall optimization objective (7). However, increasing the horizon size extends the training time. For instance, achieving a 9.1% gain in training loss convergence requires approximately 3.8 times more training time, while a 15.7% improvement demands nearly an 18-fold increase in training time. As the horizon grows, the marginal gain in convergence shrinks while the marginal cost in training time grows steeply, which is a trade-off that requires a balance.
@@ 4. Training NeRC
We train a separate NeRC for each area or route, just like other data-driven regression paradigms that tightly depend on the specific scenes in which they are applied (Mildenhall et al., 2020; Zhao et al., 2023, 2024b). Additionally, a separate NeRC model is also trained for each specific time interval of the day, as GNSS satellites (Tx) orbit the Earth periodically. In this section, we discuss how to train the NeRC pipeline. In addition to supervised learning using GNSS data labeled with ground truth locations, we propose an alternative training paradigm to unleash the power of unlabeled data.
@@ 4.1. Supervised Learning
These figures illustrate the publicly available route information from Google My Maps and the EDF cost map derived from the map information.
@@ 4.2. Unleashing the Power of Unlabeled Data
Innumerable unlabeled GNSS data are generated from mobile devices every day, but have never been fully used. As the blessings of end-to-end learning, we propose a brand new learning paradigm using Euclidean Distance Field (EDF) cost maps to unleash the power of unlabeled GNSS data. The core idea behind this innovation is to construct a differentiable map supervision by comparing the predicted trajectories with the passable routes provided by publicly available geodetic maps, like Google Maps or OpenStreetMap. This novel learning paradigm is inspired by robot motion planning (Ratliff et al., 2009; Han et al., 2019; Yang et al., 2023b), robust flight control (actual–reference trajectory comparison) (Wang et al., 2024), and map matching (Paglieroni, 1992; Liu et al., 2022).
We assume that a user travels along the passable routes in an area. As shown in 7(a), we can extract passable routes from Google Maps and construct a corresponding spatial cost function. The design principle is to ensure a lower cost value as the user’s trajectory becomes more aligned with the reference route. To this end, we propose to employ the Euclidean distance from each predicted location to its nearest reference route as the cost to supervise the learning process. This can be done through the Euclidean distance transform (EDT) of the map as illustrated by 7(b), which is referred to as the Euclidean distance field (EDF) cost map (Han et al., 2019; Yang et al., 2023b). For each spatial point on the map, EDT produces a potential value representing the distance between and its nearest point on the reference route (Felzenszwalb and Huttenlocher, 2012):
@@
The gradient descending directions of the cost map are illustrated by the white arrows in 7(b). The gradient information of the cost always pushes the location prediction toward the reference trajectory, which is a dual analogy for collision avoidance in robot motion planning (Ratliff et al., 2009; Han et al., 2019; Yang et al., 2023b).

### v2

@@ Abstract.
GNSS localization using everyday mobile devices is challenging in urban environments, as ranging errors caused by the complex propagation of satellite signals and low-quality onboard GNSS hardware are blamed for undermining positioning accuracy. Researchers have pinned their hopes on data-driven methods to regress such ranging errors from raw measurements. However, the grueling annotation of ranging errors impedes their pace. This paper presents a robust end-to-end Neural Ranging Correction (NeRC) framework, where localization-related metrics serve as the task objective for training the neural modules. Instead of seeking impractical ranging error labels, we train the neural network using ground-truth locations that are relatively easy to obtain. This functionality is supported by differentiable moving horizon location estimation (MHE) that handles a horizon of measurements for positioning and backpropagates the gradients for training. Even better, as a blessing of end-to-end learning, we propose a new training paradigm using Euclidean Distance Field (EDF) cost maps, which further alleviates the demands on labeled locations. We evaluate NeRC on public benchmarks and our collected datasets, demonstrating its distinguished improvement in positioning accuracy. We also deploy NeRC on the edge to verify its real-time performance for mobile devices.
@@ 1. Introduction
The ranging errors caused by multipath and NLOS effects are strongly dependent on the complex relationship among satellites (Tx), receivers (Rx), and the surroundings (Sun et al., 2023; Weng et al., 2024b; Sun et al., 2025). Although modeling errors in ranging measurements is challenging, they often exhibit a periodical pattern at a given site because GNSS satellites fly back approximately every 24 hours (Axelrad et al., 2005; Phan et al., 2013). For example, we collected GNSS data using an Android phone at the same site and time across multiple days and visualized ranging errors in 1(b). The color of each satellite’s trajectory indicates the value of ranging errors, estimated from ground-truth user locations after removing all known error sources, including satellite clock offsets, atmospheric delays, and relativistic effects (Weng et al., 2024b). 1(b) shows that ranging errors at a specific site and time exhibit a similar pattern.
The context-dependence and predictability of ranging errors have propelled the community away from traditional model-based approaches (Axelrad et al., 2005; Liu et al., 2018b; Bai et al., 2020) toward data-driven methods (Sun et al., 2020, 2021; Zhang et al., 2021; Sun et al., 2023; Weng et al., 2024b; Sun et al., 2025). However, obtaining accurate labels for ranging errors—critical for training neural models—is both difficult and impractical. As illustrated by Figure 2a, previous studies have used various sophisticated algorithms to estimate these errors based on ground truth locations provided by high-performance specialized positioning systems (Zhang et al., 2021; Sun et al., 2023; Weng et al., 2024b; Sun et al., 2025). With the derived ranging errors, an intermediate loss can be computed to train a neural network for ranging correction. Yet, manually calculating ranging errors is inherently imprecise due to the unobservable user clock offsets and stochastic noise (Weng et al., 2024b). This raises the question: can we instead utilize these location labels directly, given that our ultimate goal is localization? If so, we could bypass the need to compute intermediate losses using ranging error labels and instead train neural modules end-to-end with the final positioning task loss.
@@
Can we integrate the two learning paradigms (shown in Figure 2a and 2b) to harness the powerful regression capabilities of neural networks while preserving our interpretable knowledge of GNSS localization? The key ingredient to enable such an end-to-end learning framework should be a differentiable positioning engine, as illustrated in Figure 2c. This engine leverages outputs from an upstream neural model to compute locations while backpropagating gradients from final positioning tasks to refine the upstream neural model. This structured learning paradigm, widely adopted in fields such as robotics (Wang et al., 2024; Yang et al., 2023; Wang et al., 2025), autonomous driving (Huang et al., 2024), and 3D vision (Mildenhall et al., 2020; Chen et al., 2025), has recently been extended to GNSS positioning (Xu et al., 2023a; Weng et al., 2024c; Xu and Hsu, 2024; Hu et al., 2025).
Among the most relevant studies, E2E-PrNet and pyrtklib train upstream neural networks to predict pseudorange errors through a differentiable Weighted Least Squares (WLS) localization engine (Weng et al., 2024c; Hu et al., 2025). While these methods improve positioning in rural (Weng et al., 2024c) or lightly urban areas (Hu et al., 2025), they struggle in deep urban environments, as shown in Figure 3. The main reason is that the WLS engine lacks robustness to the severe noise present in such settings, where multipath and NLOS propagation typically cause low carrier-to-noise density ratios () (Weng and Ling, 2023). Furthermore, these methods rely solely on fully labeled 3D location data, limiting their practical deployment since obtaining accurate ground-truth locations is both challenging and costly (Fu et al., 2020).
@@ 2. NeRC Design
NeRC is a data-driven framework guided by well-established physical and geometric priors. It handles a horizon of raw measurements for neural ranging error regression and robust location estimation. Within it, the neural module is trained with positioning-related metrics. This end-to-end learning pipeline—from raw measurements to locations—comprises two core modules:
@@ 2.1. Upstream Network for Ranging Correction
Figure 4 illustrates that one input tensor sample has two dimensions corresponding to satellites and features. The MLP operates in a satellite-wise manner—analogous to the position-wise MLPs in the transformer architecture—processing information from all satellites in parallel. A visibility mask handles the time-varying set of visible satellites by zeroing the outputs of invisible ones, analogous to the transformer’s masked softmax. For clarity, Figure 4 presents an unfolded view of the input tensors from time to time . In reality, the NeRC pipeline uses a single MLP that processes the entire time window in parallel. Including the time dimension and the batch dimension, the resulting four-dimensional input tensor has the shape , where is the batch size, is the total number of satellites, and is the number of input features. Thus, the MLP can be represented by
where denotes the 4D input tensor with a horizon size of at time , is the corresponding pseudorange error predictions, and represents all learnable parameters in the neural network.
@@ 2.2.2. Moving Horizon Estimation
where denotes an estimate of an unknown at time given measurements up to time . and represent the covariance matrices of process and measurement noise, respectively. The inverse of them is used to weight the state transition and measurement costs. is the covariance matrix of the state estimation , representing prior knowledge of our confidence in the initial user state for the current horizon . It also represents the “inertia” of the system—a larger increases the contribution of the prior state cost (arrival cost) in the overall objective, thereby drawing the state estimates in the current horizon more consistent with historical trajectories. Assuming the system statistics in (7) are Gaussian, MHE formulated by (7) is equivalent to the Maximum A Posteriori (MAP) estimation (Jazwinski, 1970; Weng et al., 2025):
@@ 2.2.3. Differentiable Formulation of MHE
where is treated as prior knowledge and is initialized as a diagonal matrix with all diagonal elements tuned empirically. is the Jacobian matrix of the nonlinear function :
@@
Figure 5 illustrates the principle of the differentiable MHE. By substituting (9) (13) into (7), we obtain the state transition cost, measurement cost, and prior state cost. The objective layer aggregates the three types of costs weighted by their corresponding covariance matrices, which is then solved by a general Differentiable Nonlinear Least Squares (DNLS) solver. In our implementation, we employ the Theseus library built for Pytorch (Pineda et al., 2022). The optimizer is initialized with an approximate trajectory guess, denoted by , for each horizon. We leverage the historical state estimates for initialization (Weng et al., 2025). The lower half of this diagram depicts the updating process of the prior state covariance matrix using (14), where only the latest covariance matrix gets updated at time . The updated covariance matrices serves the next horizons.
@@ 3.2. Forward profiling
6(a) indicates that MHE-w/-AC maintains a lead over other algorithms as the horizon size is small. By contrast, MHE-w/o-AC leverages measurements within the current horizon and disregards prior information, which significantly degrades its performance in short horizons. Increasing the horizon size can narrow down the performance gap between MHE-w/-AC and MHE-w/o-AC. Once the horizon extends beyond a certain threshold—approximately 65 in this example—MHE-w/o-AC begins to outperform MHE-w/-AC. At that point, the window is sufficiently long to retain informative measurements for current-state estimation, while older data outside the window degrades performance (Ling and Lim, 1996, 1999; Weng et al., 2025). Continuously increasing the horizon size does not always lead to improved performance. A typical performance curve of MHE-w/o-AC initially decreases to a minimum, then gradually converges back toward that of MHE-w/-AC (Rao and Rawlings, 2002; Weng et al., 2025). In this dataset, we observe that the turning point occurs at a horizon length of approximately 450. The filtering-based MHE-F exhibits the same performance regardless of the chosen horizon size (Muske et al., 1993; Weng et al., 2025).
Regarding the computation time of each method, 6(b) shows that the forward time increases as the horizon size enlarges. Among the methods, MHE-F demonstrates the fastest performance, benefiting from its recursive filtering structure. In contrast, the optimization-based approaches are naturally slower, as they solve a nonlinear least squares problem at each time step. Notably, MHE-w/-AC incurs slightly more computation than MHE-w/o-AC for the same horizon size, due to the additional optimization for the arrival cost. However, MHE-w/-AC can achieve superior positioning accuracy with a smaller horizon, thereby taking less computational time to reach similar performance levels.
Summary: MHE can offer advanced positioning capabilities at a competitive cost with appropriate configurations of its horizon size and arrival cost. Increasing the horizon size can enhance the accuracy of MHE-w/o-AC, but the marginal gains over MHE-w/-AC with a smaller horizon come with a substantial rise in computational time—a 2.35% improvement in accuracy comes at the expense of approximately 2.4 times the computation time, as shown in 6(a) and 6(b). This trade-off is critical to consider when deploying MHE on resource-constrained platforms such as mobile devices. Additional improvement in MHE is outside the scope of this paper, but can be expected by approximating the arrival cost more accurately (Rao and Rawlings, 2000; Tenny and Rawlings, 2002; Rao et al., 2003; Fiedler et al., 2020) and incorporating constraints on the system (Rao et al., 2001; Haseltine and Rawlings, 2005; Wen et al., 2021; Suzuki, 2023; Xia et al., 2024).
@@ 3.3. Backward profiling
This section examines the training convergence of NeRC, an upstream MLP equipped with MHE. 6(c) gives the training loss (3D location loss) after one training epoch for various horizon sizes.
Increasing the horizon size improves training convergence across all estimators. While MHE-w/o-AC exhibits poor convergence at small horizons, it surpasses the other two methods in convergence speed once the horizon exceeds 20, as illustrated in 6(c). It is the arrival cost item in (7)—the system’s ‘inertia’—that decelerates the backward convergence of MHE-F and MHE-w/-AC. Interestingly, it is the same item that enables MHE to achieve satisfactory forward positioning accuracy even with small horizon sizes.
6(d) shows the backward training time per batch of the three methods. Just aligned with the forward time, MHE-F exhibits the lowest backward computation time due to its filtering structure.
Summary: The backward training convergence can be improved by expanding the receding horizon, which is equivalent to reducing the contribution of the arrival cost in the overall optimization objective (7). However, increasing the horizon size extends the training time. For instance, achieving a 9.1% gain in training loss convergence requires approximately 3.8 times more training time, while a 15.7% improvement demands nearly an 18-fold increase in training time. As the horizon grows, the marginal gain in convergence shrinks while the cost in training time grows steeply, which is a trade-off that requires a balance.
@@ 4. Training NeRC
We train a separate NeRC for each area or route, just like other data-driven regression paradigms that tightly depend on the specific scenes in which they are applied (Mildenhall et al., 2020; Zhao et al., 2023, 2024b). Additionally, separate NeRC models should be trained for specific time intervals of a day, as GNSS satellites (Tx) orbit the Earth periodically. Since pseudorange errors are also closely related to device variations (Weng et al., 2024b), NeRC needs to be retrained for different mobile phones. We specify how to train NeRC here. In addition to supervised learning using data labeled with true locations, we propose an alternative training paradigm to unleash the power of unlabeled data.
@@ 4.2. Unleashing the Power of Unlabeled Data
Innumerable unlabeled GNSS data are generated from mobile devices every day, but have never been fully used. As the blessings of end-to-end learning, we propose a brand new learning paradigm using Euclidean Distance Field (EDF) cost maps to unleash the power of unlabeled GNSS data. The core idea behind this innovation is to construct a differentiable map supervision by comparing the predicted trajectories with the passable routes provided by publicly available geodetic maps, like Google Maps or OpenStreetMap. This novel learning paradigm is inspired by robot motion planning (Ratliff et al., 2009; Han et al., 2019; Yang et al., 2023), robust flight control (actual–reference trajectory comparison) (Wang et al., 2024), and map matching (Paglieroni, 1992; Liu et al., 2022).
We assume that a user travels along the passable routes in an area. As shown in 7(a), we can extract passable routes from Google Maps and construct a corresponding spatial cost function. The design principle is to ensure a lower cost value as the user’s trajectory becomes more aligned with the reference route. To this end, we propose to employ the Euclidean distance from each predicted location to its nearest reference route as the cost to supervise the learning process. This can be done through the Euclidean distance transform (EDT) of the map as illustrated by 7(b), which is referred to as the Euclidean distance field (EDF) cost map (Han et al., 2019; Yang et al., 2023). For each spatial point on the map, EDT produces a potential value representing the distance between and its nearest point on the reference route (Felzenszwalb and Huttenlocher, 2012):
@@
The gradient descending directions of the cost map are illustrated by white arrows in 7(b) and always push location predictions toward reference trajectories, which is a dual analogy for collision avoidance in robot motion planning (Ratliff et al., 2009; Han et al., 2019; Yang et al., 2023). Even when data collection routes slightly deviate from reference trajectories, considerable improvement can still be achieved, since GNSS positioning errors in urban areas often span hundreds of meters, exceeding street widths. Note also that EDF cost maps serve here only as loss functions for training with unlabeled data, not as optimization constraints. And only one EDF cost map is needed to train NeRC for a given region across different times of day. The inference process does not require any map information.